# Technologies for Modulation of Visible and Near-UV Light and their Applications

### v1

@@ III Integrated Photonics
Here we provide an overview of recent advances in active visible photonic integrated circuits (PICs). Visible PICs have applications in low linewidth lasers, optical frequency combs, and sensing. These applications have been covered in a previous review [160] and are not discussed here. Instead, we focus on devices that provide the amplitude and phase control needed for the applications laid out in Sec. I. In particular, we review four nonlinearities used for integrated visible light modulation - thermo-optic tuning (Sec. III.1), electro-optic tuning (Sec. III.2), piezo-optic tuning (Sec. III.3), and liquid-crystal-based modulators (Sec. III.4).
@@
[…]
Integrated modulators can also be categorized by their functionality: phase-only, amplitude-only, or phase and amplitude modulators. In this review we do not distinguish between these categories as the physical mechanism is generally the same and device design can convert between types as needed for a given application. Additionally, most integrated solutions focus on single-channel modulation but hybrid systems or active channel multiplexing (e.g. optical phased arrays [166]) can be used to control the spatial distribution in the far field.
III.1 Integrated Thermo-Optic Modulators
@@
Thermo-optic (TO) tuning provides a straight forward way to change the real part of a waveguide’s effective refractive index, and is compatible with all materials. In a TO modulator, a resistive heater changes the temperature, and thus refractive index, of a photonic device. Due to its simplicity and versatility, TO modulation has been extensively employed to provide compact phase and amplitude control.
[…]
@@
SiN’s fairly large bandgap and mature fabrication techniques have enabled integrated active devices with operating wavelengths below 500 nm, necessary for all three of the applications discussed Sec. I, including devices with 50 kHz bandwidth for applications in neural stimulation [171] and systems with high channel counts for visible-light optical phased arrays [172].
@@
In general, TO modulators suffer from thermal crosstalk and an innate tradeoff between power consumption and switching bandwidth. Standard devices are hundreds of micrometers in length and require tens of mW for a $\pi$ phase shift. However, thermal and optical engineering can improve these specifications [173]. For example, power consumption below 1 mW for $\pi$ phase shift ($W_{\pi}$) has been demonstrated using suspended devices and multi-pass waveguides, though with switching times of only $\sim 1$ kHz [174, 175]. Similarly, microresonator based TO modulators leverage resonant enhancement to reduce power consumption by a factor proportional to the resonator’s finesse. This technology has been used to demonstrate devices with 3dB bandwidths of nearly 300 kHz with $W_{\pi}\sim 1$ mW [176].
III.2 Integrated Electro-Optic Modulators
Here we review integrated versions of the bulk electro-optic modulators discussed in Sec. II.
III.2.1 Lithium Nioabte
@@
As discussed in Sec. II, bulk LiN modulators are the workhorse of commercial light modulation but integrated LiN devices have been hampered by the difficulty of creating and patterning thin films. However, developments in commercial production of thin film LiN-on-insulator at the wafer scale [177] and fabrication methods [178] have spurred a recent surge of integrated LiN photonic devices across the optical spectrum [179, 180, 181, 182]. In particular, three recent publications have demonstrated fast modulation in the near-infrared. In [183], a rib waveguide is used to reduce loss due to sidewall roughness and defects at the LiN-oxide cladding interface while retaining single mode propagation at 635 nm and low bending loss. With these devices they demonstrated high speed switching at 850 nm with a 3 dB cutoff at 10 GHz and a $V_{\pi}=8$ V for a 2 mm device (1.6 V-cm). Device improvements enabled sub-V-cm, 35 GHz bandwidth switching at 780 nm [184]. In [185] Mg-doped LiN on a sapphire substrate is used to reduce rf losses, and device design and improved fabrication enabled reduced optical losses, device size, and voltage requirements. These improvements enable modulation of 780 nm light at 2.7 GHz and the CMOS-compatible voltage of 4.2 V for a 3 mm device (1.4 V-cm). Finally, [186] uses full system-level optimization to demonstrate a multi-channel architecture with modulation of 780 nm at 7 GHz with 2.2 V control voltage for a 3 mm device (0.66 V-cm) and larger than 20 dB extinction. Recently, full-spectrum visible modulation was demonstrated at 450, 520, and 630 nm with sub V-cm control voltages, however with only modest extinction ratios [187].
@@
All of these results boast fairly low switching voltages and large switching bandwidths. This comes at the expense of fairly large device size (mm-scale). Due to the large bandgap of LiN, there is no fundamental reason that these results cannot be extended to lower wavelengths although the fabrication will be more difficult and etch roughness will lead to more loss.
@@
Unfortunately, LiN integrated photonics are plagued by phase drifts presumably due to the photo-refractive effect [188, 186, 189] in which the guided optical fields themselves change the local refractive index. This is presumably due to fluctuations of trapped charges, although the exact mechanism is still unknown. This causes slow drifts in the operating parameters of modulators, especially at high optical powers [190, 191]. Work has been done to mitigate this effect by removing the cladding [189] or through active stabilization [186]. Additionally, the pyro-electric coefficient in LiN can limit fabrication and is especially a concern for cryogenic integrated photonics [192, 193] as it can unpredictably change the electrical and optical properties during the cool down from room temperature. [194]. Although large-scale systems have been demonstrated despite these issues [186], more research will help in understanding and mitigating these effects.
III.2.2 Aluminum Nitride
@@ III.2.3 Emerging Materials
We have focused on LiN and AlN due to their prevalence as integrated optics platforms, but other materials are emerging as possible solutions for fast EO modulation in the visible spectrum. For instance, Barium Titanate (BTO) has a large bandgap of 3.2 eV and a strong EO coefficient of up to 1300 pm/V for unclamped films [200, 201]. BTO films deposited on Magnesium Oxide substrates have been shown to have low loss at 633 nm and large ridge-waveguide EO modulators were fabricated providing up to 1 MHz modulation at 633 nm [202]. Graphene can also be used for EO modulation with incredibly compact form factors (10s of $\mu$m) and fast modulation (10s of GHz) for single devices [203, 204], and even on the wafer scale [205]. Although these results focused on IR modulation, similar techniques could be translated to modulation of visible light as required for the applications discussed in this review.
III.3 Integrated Strain-Tuning-Based Modulators
@@
In this section we discuss devices which use physical deformation of photonic structures for visible light modulation. This effect takes many forms depending on the material platform. In all instances a change in the local refractive index or device geometry due to pressure or strain is used to modulate an incident laser.
@@
As discussed in Sec. II.2.1, AOMs, which make use of Brillouin scattering to modulate light, are a commercial mainstay for visible light modulation. While these commercial products are generally based on bulk crystals, the first demonstrations of acoustic-optic devices focused on modulation of light in a 2D guided mode with surface acoustic waves. In a seminal publication in 1970, 633 nm light was guided in a thin glass film deposited on the surface of a $\alpha$-quartz crystal providing modulation efficiencies of up to 66% with 0.18 W of acoustic power [206]. In 1974, a device employing 2 tilted surface acoustic waves in an optical guiding layer of LiN enabled high bandwidth (200 MHz) modulation of 633 nm light with a diffraction efficiency of 50% at 0.2 W rf power [207] and later device improvements enabled amplitude and frequency modulation without position modulation, imperative for applications which call for fiber and waveguide coupling [208]. Because these guided-wave acoustic-optic devices confine both optical and acoustic waves in a 2D planar structures, the light-sound interaction is enhanced, improving efficiency, bandwidth, and power consumption.
@@
More recently, an integrated AOM built on LiN was demonstrated in the infrared for LiDAR systems [209]. In this demonstration, the acoustic wave and optical wave are anti-parallel – in contrast to most commercial devices and previous planar devices where the acoustic wave propagates perpendicularly to the optical field. With the anti-parallel geometry, the light is directly diffracted out of the chip. In this device, the authors achieve a switching bandwidth of 0.7 MHz, likely limited by electronics. As LiN boasts a large bandgap, this technology could be translated to the visible, with direct applications in atom control [210] which, as discussed in Sec. I relies on 2D beam positioning to move the atoms for different operations. Patterning waveguides to achieve full confinement could further improve optical performance. In a recent demonstration, on-chip acoustic modulation was used to frequency shift and deflect an integrated optical field, an important step towards a fully integrated device [211]. Again, although demonstrated at 1597 nm, the LiN platform should allow for direct translation to visible modulation.
@@
In contrast to the above works which use an RF field to induce a controllable grating in a piezo-optic material, more recent works directly deform a nano-photonic device. This idea was initially discussed in [212] which considered modulation of a buried SiN waveguide. With a high piezo-electric nonlinearity, AlN is an obvious material to take advantage of this effect. However, as discussed above, achieving low-loss propagation in AlN has proven difficult. To get around this limitation a set of recent works use a hybrid approach of guiding light in SiN waveguides with the physical geometry or refractive index modulated with actively strained AlN.
@@
A recent set of publications demonstrated integrated modulation of SiN devices at visible wavelengths using AlN as the piezoelectric material. In [216], a 40 $\mu$m diameter SiN ring resonator is modulated at 250 MHz with 0.125 mW of electrical power at room temperature, and even lower power consumption when operated at 7 K. Additionally, a waveguide-based Mach-Zender modulator (MZM) is modulated in 750 ns with 20 nW power dissipation. However, these MZMs suffered from large insertion loss and high voltage requirements, leading to a 175 VdB voltage-loss product and limiting the utility in large scale systems. In a follow-up work, an undercut of the modulated structures improved the voltage requirements and improved loss to achieve a 0.8 VdB voltage-loss product [217], although with a modulation speed limited to 4 MHz by the mechanical resonance. Finally, in [218] simultaneous optimization of all optical, electrical, and mechanical components results in devices with 50 MHz modulation with a power dissipation of only 200 $\mu$W at 1 MHz modulation. A variation on these devices has been characterized with atomic quantum control in mind [219], focusing on systems-scale requirements such as repeatability, extinction, and scalability.
@@
Figure 4: (a) Qualitative depiction of the capabilities of integrated modulators using thermal, electric, mechanical, and liquid-crystal modulation. (b) Wavelength vs modulation speed for all technologies discussed in this review.
III.4 Integrated Liquid-Crystal-Based Modulators
@@
Compared to integrated thermo-optic, electro-optic, and piezo-optomechanical modulators, integrated liquid-crystal-based (LC-based) modulators enable very compact form factors and power-efficient operation, at the expense of slower modulation speeds (typically sub-KHz) [224, 225]. These LC-based modulators leverage the large birefringence and electrically-tunable refractive index of LC media by integrating LC regions into integrated-photonics platforms.
@@
These LC-based modulators typically consist of a waveguide that guides the light near a trench filled with the LC medium. The optical mode of the waveguide then overlaps and interacts with the integrated LC medium. Consequently, the effective refractive index and profile of the optical mode are influenced by the refractive index of the LC region. The refractive index of the LC medium seen by the optical mode depends on the orientation of the LC molecules with respect to the direction of light propagating in the waveguide. As discussed above (Sec. III.4), the orientation of the LC molecules can be tuned by applying a voltage across the LC region using integrated electrodes, since LC molecules align to an external electric field. As the LC molecules rotate, the refractive index of the LC medium changes, resulting in a change in the effective index and profile of the optical mode.
@@
Various commercially-available LC compounds can be used for these modulators, with the compound used impacting parameters such as the birefringence at a given wavelength, the operating temperature range, and the modulation speed. For example, a commonly-utilized LC compound is 5CB, which has ordinary and extraordinary refractive indices of 1.53 and 1.7, respectively, at red wavelengths and a nematic-state operating temperature range of 18 to ${}^{\circ}$.
@@
Integrated LC-based demonstrations have largely focused on operation at infrared wavelengths [224, 225, 226, 227, 228, 229, 230, 231, 232, 233]. Recently, integrated LC-based modulators have been shown at visible wavelengths by integrating nematic LC media into silicon-nitride-based platforms [234, 235, 236, 237, 238], enabling compact phase and amplitude modulation at a red operating wavelength. For example, an integrated LC-based phase modulator achieved $41\pi$ phase shift within ±$\pm 2.4$ V for a 500-$\upmu$m-long modulator, which means that a $2\pi$ phase shifter would need to be only 24.4 $\upmu$m long (corresponding to a 0.003 V-cm modulation requirement) [234, 235]. An integrated LC-based amplitude modulator leveraging the LC birefringence to tune the coupling coefficient between two waveguides, achieved amplitude modulation with a device length of only 14 $\upmu$[237, 236].
@@
Due to the large birefringence afforded by LC media, these integrated LC-based devices are one to two orders of magnitude more compact than comparable thermo-optic, electro-optic, or piezo-optomechanical modulators at visible wavelengths, making the particularly suited to applications [239, 240, 7, 101, 8], including augmented-reality displays (as discussed in Sec. I). For adoption in quantum and biological applications, the modulation speed of these integrated LC-based modulators will need to be improved.
III.5 Summary

### v2

@@ III Integrated Photonics
Here we provide an overview of recent advances in active visible photonic integrated circuits (PICs). Visible PICs have applications in low linewidth lasers, optical frequency combs, and sensing. These applications have been covered in a previous review [172] and are not discussed here. Instead, we focus on devices that provide the amplitude and phase control needed for the applications laid out in Sec. I. In particular, we review four nonlinearities used for integrated visible light modulation - thermo-optic tuning (Sec. III.2), electro-optic tuning (Sec. III.3), piezo-optic tuning (Sec. III.4), and liquid-crystal-based modulators (Sec. III.5).
@@
In general, optical loss in a PIC is due to a combination of (i) material absorption; (ii) scattering due to device fabrication imperfections; and (iii) scattering due to material defects. Loss due to all three of these sources increase at shorter wavelengths: absorption increases near a material’s band edge and Rayleigh scattering due to surface roughness or material imperfections scales as $\lambda^{-4}$ [173, 174]. Nonetheless, low-loss propagation is possible, especially with the field’s continual advances in material growth and processing. In this section we will discuss active PIC devices with a guiding layer of Silicon Nitride, Aluminum Nitride, or Lithium Niobate each of which has its own benefits and drawbacks. Other wide bandgap materials such as Aluminum Oxide (Al2O3[175], and diamond [176] can support visible PICs, though we do not discuss them here as there have not been demonstrations of active modulation with these material systems.
III.1 Principle of Operation
Figure 6: (a) Method for phase modulation - an external signal changes the index of refraction of portion of a waveguide, adding a controllable phase to the light through the waveguide. (b) The phase modulation creates amplitlude modulation when embedded in a Mach Zender interferometer.
Before discussing particular results, we briefly lay out the basic principle of operation for integrated modulators. In all integrated modulators, an external control is used to modulate the refractive index of a material, which in turn changes the effective refractive index of a guided mode, as illustrated in Fig. 6. This enables control of the phase of the optical mode. Embedding such a device in an interferometer enables amplitude as well as phase control. Device design can optimize parameters of interest such as modulator speed, size, or power consumption. However, in this review we focus on the modulation mechanism and materials that make visible light modulation possible and not on advances in device geometry.
III.2 Integrated Thermo-Optic Modulators
@@
Thermo-optic (TO) tuning provides a straight forward way to change the real part of a waveguide’s effective refractive index, and is compatible with all materials. As seen in Fig. 7, in a TO modulator, a resistive heater changes the temperature, and thus refractive index, of a photonic device. Due to its simplicity and versatility, TO modulation has been extensively employed to provide compact phase and amplitude control.
Figure 7: (a) Simplified cross-sectional diagram of an example electro-optic phase modulator, consisting of a waveguide with control electrodes on each side (not to scale). (b) Simplified top-view schematic of the device.
TO modulators have been used extensively with Silicon Nitride (SiNx) photonic devices. SiN boasts record-low propagation loss [177] but lacks even-order nonlinearities due to its centrosymmetric molecular structure. Stoichiometric Si3N4 deposited by low pressure chemical vapor deposition (LPCVD) is transparent above about 400 nm. High-temperature annealing removes hydrogen and reduces defect absorption. Si3N4 is also CMOS-compatible and structures can be fabricated on the wafer-scale. The vast field of SiN integrated photonics has been reviewed extensively [178, 179, 180]. Many TO devices are used for static control. Here, we focus on TO modulation results providing switching speeds at or above 1 kHz.
@@
SiN’s fairly large bandgap and mature fabrication techniques have enabled integrated active devices with operating wavelengths below 500 nm, necessary for all three of the applications discussed in Sec. I. Devices with 50 kHz switching bandwidth have been demonstrated for applications in neural stimulation [127]. Additionally, mature dSiN fabrication enables system-scale fabrication with high channel counts for visible-light optical phased arrays [181].
@@
TO modulators suffer from thermal crosstalk and an innate tradeoff between power consumption and switching bandwidth. Standard devices are hundreds of micrometers in length and require tens of mW for a $\pi$ phase shift. However, thermal and optical engineering can improve these specifications [182]. For example, power consumption below 1 mW for $\pi$ phase shift ($W_{\pi}$) has been demonstrated using suspended devices and multi-pass waveguides, though with switching times of only $\sim 1$ kHz [183, 184]. Similarly, microresonator based TO modulators leverage resonant enhancement to reduce power consumption by a factor proportional to the resonator’s finesse. This technology has been used to demonstrate devices with 3dB bandwidths of nearly 300 kHz with $W_{\pi}\sim 1$ mW [185]. In general, Si3N4 TO modulators provide a relatively mature solution for applications which can tolerate lower speeds and don’t have strict power or temperature constraints.
III.3 Integrated Electro-Optic Modulators
Here we review integrated versions of the bulk electro-optic modulators discussed in Sec. II. Independently of the material used, an electro-optic modulator comprises a waveguide and two electrodes. An common geometry for a phase-only electro-optic modulator is seen in Fig. 8. A voltage applied between the electrodes creates an electric field across the waveguide which creates a local change in the index of refraction of a material with an electro-optic nonlinearity.
Figure 8: (a) Simplified cross-sectional diagram of an example electro-optic phase modulator, consisting of a waveguide with control electrodes on each side (not to scale). (b) Simplified top-view schematic of the device.
Figure 9: (a) Qualitative depiction of the capabilities of integrated modulators using thermal, electric, mechanical, and liquid-crystal modulation. (b) Wavelength vs modulation speed for all technologies discussed in this review. Lines represent bulk technologies discussed in Sec. II.
III.3.1 Lithium Nioabte
@@
As discussed in Sec. II, bulk LiNbO3 modulators are the workhorse of commercial light modulation but integrated LiNbO3 devices have until recently been elusive due to the difficulty of creating and patterning thin films. However, developments in commercial production of thin film LiNbO3 -on-insulator at the wafer scale [186] and fabrication methods [187] have spurred a recent surge of integrated photonic devices across the optical spectrum [188, 189, 190, 191]. In particular, three recent publications have demonstrated fast modulation in the near-infrared. In [192], a rib waveguide is used to reduce loss due to sidewall roughness and defects at the LiNbO3 -oxide cladding interface while retaining single mode propagation at 635 nm and low bending loss. With these devices they demonstrated high speed switching at 850 nm with a 3 dB cutoff at 10 GHz and a $V_{\pi}=8$ V for a 2 mm device (1.6 V-cm). Further device improvements enabled sub-V-cm, 35 GHz bandwidth switching at 780 nm [193]. In [194] Mg-doped LiNbO3 on a sapphire substrate is used to reduce rf losses, and device design and improved fabrication enabled reduced optical losses, device size, and voltage requirements. These improvements enable modulation of 780 nm light at 2.7 GHz and the CMOS-compatible voltage of 4.2 V for a 3 mm device (1.4 V-cm). Finally, [195] uses full system-level optimization to demonstrate a multi-channel architecture with modulation of 780 nm at 7 GHz with 2.2 V control voltage for a 3 mm device (0.66 V-cm) and larger than 20 dB extinction. Recently, full-spectrum visible modulation was demonstrated at 450, 520, and 630 nm with sub V-cm control voltages [196].
@@
All of these results boast fairly low switching voltages and large switching bandwidths. This comes at the expense of fairly large device size (mm-scale). Due to the large bandgap of LiNbO3 , there is no fundamental reason that these results cannot be extended to lower wavelengths although the fabrication will be more difficult and etch roughness will lead to more loss.
@@
Unfortunately, LiNbO3 integrated photonics are further plagued by phase drifts, presumably due to the photo-refractive effect [197, 195, 198] in which the guided optical fields themselves change the local refractive index. This is presumably due to fluctuations of trapped charges, although the exact mechanism is still unknown. This causes slow drifts in the operating parameters of modulators, especially at high optical powers [199, 200]. Work has been done to mitigate this effect by removing the cladding [198] or through active stabilization [195]. Additionally, the pyro-electric coefficient in LiNbO3 can limit fabrication and is especially a concern for cryogenic integrated photonics [201, 202] as it can unpredictably change the electrical and optical properties during the cool down from room temperature [203]. Although large-scale systems have been demonstrated despite these issues [195], more research is necessary to understand and mitigate these effects.
III.3.2 Aluminum Nitride
@@ III.3.3 Emerging Materials
We have focused on LiNbO3 and AlN due to their prevalence as integrated optics platforms, but other materials are emerging as possible solutions for fast EO modulation in the visible spectrum. For instance, Barium Titanate (BTO) has a large bandgap of 3.2 eV and a strong EO coefficient of up to 1300 pm/V for unclamped films [209, 210]. BTO films deposited on Magnesium Oxide substrates have been shown to have low loss at 633 nm and large ridge-waveguide EO modulators were fabricated providing up to 1 MHz modulation at 633 nm [211]. Graphene can also be used for EO modulation with incredibly compact form factors (10s of $\mu$m) and fast modulation (10s of GHz) for single devices [212, 213], and even on the wafer scale [214]. Although these results focused on infrared modulation, similar techniques could be translated to modulation of visible light as required for the applications discussed in this review. Finally, Silicon Carbide (SiC), a wide bandgap material, is emerging as a platform for integrated photonics [215], especially as it is CMOS-compatible. Recent improvements in fabrication quality enabled GHz modulation at infrared wavelengths in 3C-SiC [216] and further fabrication improvements could enable visible modulation as well.
III.4 Integrated Strain-Tuning-Based Modulators
@@
In this section we discuss devices which use physical deformation of photonic structures for visible light modulation. This effect takes many forms depending on the material platform. In all instances a change in the local refractive index or device geometry due to strain is used to modulate light routed through a waveguide.
@@
As discussed in Sec. II.2.1, AOMs, which make use of Brillouin scattering to modulate light, are a commercial mainstay for visible light modulation. While these commercial products are generally based on bulk crystals, the first demonstrations of acoustic-optic devices focused on modulation of light in a 2D guided mode with surface acoustic waves. In a seminal publication in 1970, 633 nm light was guided in a thin glass film deposited on the surface of a $\alpha$-quartz crystal providing modulation efficiencies of up to 66% with 0.18 W of acoustic power [217]. In 1974, a device employing 2 tilted surface acoustic waves in an optical guiding layer of LiNbO3 enabled high bandwidth (200 MHz) modulation of 633 nm light with a diffraction efficiency of 50% at 0.2 W rf power [218] and later device improvements enabled amplitude and frequency modulation without position modulation, imperative for applications which call for fiber and waveguide coupling [219]. Because these guided-wave acoustic-optic devices confine both optical and acoustic waves in a 2D planar structures, the light-sound interaction is enhanced, improving efficiency, bandwidth, and power consumption.
@@
More recently, an integrated AOM built on LiNbO3 was demonstrated in the infrared for LiDAR systems [220]. In this demonstration, the acoustic wave and optical wave are anti-parallel – in contrast to most commercial devices and previous planar devices where the acoustic wave propagates perpendicularly to the optical field. With the anti-parallel geometry, the light is directly diffracted out of the chip. In this device, the authors achieve a switching bandwidth of 0.7 MHz, likely limited by electronics. This technology could conceivably be translated to the visible, with direct applications in atom control [221] which, as discussed in Sec. I relies on 2D beam positioning to move the atoms for different operations. Patterning waveguides to achieve full confinement could further improve optical performance. In a recent demonstration, on-chip acoustic modulation was used to frequency shift and deflect an integrated optical field, an important step towards a fully integrated device [222]. Again, although demonstrated at 1597 nm, the LiNbO3 platform should allow for direct translation to visible modulation.
@@
In contrast to the above works which use an RF field to induce a controllable grating in a piezo-optic material, more recent works directly deform a nano-photonic device. This idea was initially discussed in [223] which considered modulation of a buried SiN waveguide.
@@
A recent set of publications demonstrated integrated modulation of SiN devices at visible wavelengths using AlN as the piezoelectric material. In [227], a 40 $\mu$m diameter SiN ring resonator is modulated at 250 MHz with 0.125 mW of electrical power at room temperature, and even lower power consumption when operated at 7 K. Additionally, a waveguide-based Mach-Zender modulator (MZM) is modulated in 750 ns with 20 nW power dissipation. However, these MZMs suffered from large insertion loss and high voltage requirements, leading to a 175 VdB voltage-loss product and limiting the utility in large scale systems. In a follow-up work, an undercut of the modulated structures improved the voltage requirements and improved loss to achieve a 0.8 VdB voltage-loss product [228], although with a modulation speed limited to 4 MHz by the mechanical resonance. In [229] simultaneous optimization of all optical, electrical, and mechanical components results in devices with 50 MHz modulation with a power dissipation of only 200 $\mu$W at 1 MHz modulation. Finally, resonant modulators provide a smaller footprint with a larger modulation bandwidth (172 MHz) and high extinction ratio ($>$ 30 dB) [230]. These devices are the first designed and tested with the requirements for atomic quantum control, considering repeatability and scalability as well as device-specific parameters.
@@
III.5 Integrated Liquid-Crystal-Based Modulators
@@
Compared to integrated thermo-optic, electro-optic, and piezo-optomechanical modulators, integrated liquid-crystal-based (LC-based) modulators enable very compact form factors and power-efficient operation, at the expense of slower modulation speeds (typically sub-kHz) [235, 236]. These LC-based modulators leverage the large birefringence and electrically-tunable refractive index of LC media by integrating LC into integrated-photonics platforms.
@@
As shown in Fig. 10, these LC-based modulators typically consist of a waveguide that guides the light beneath a trench filled with LC medium, which enables the optical mode of the waveguide to overlap and interact with the integrated LC medium. Consequently, the effective refractive index and profile of the optical mode are influenced by the refractive index of the LC region. The refractive index of the LC medium seen by the optical mode depends on the orientation of the LC molecules with respect to the direction of light propagating in the waveguide. The orientation of the LC molecules can be tuned by applying a voltage across the LC region using integrated electrodes, since LC molecules align to an external electric field (as discussed above in Sec. III.5). As the LC molecules rotate, the refractive index of the LC medium changes, resulting in a change in the effective index and profile of the optical mode.
Figure 10: (a) Simplified cross-sectional diagram of an example integrated LC-based phase modulator, consisting of a silicon-nitride waveguide recessed within silicon dioxide below an LC-filled trench with integrated electrodes on each side (not to scale). (b) Simplified top-view schematic of the device when no voltage is applied across the liquid-crystal region, resulting in parallel alignment of the LC molecules with respect to the waveguide (not to scale). (c) Simplified top-view schematic of the device when the maximum voltage is applied across the LC region, resulting in complete rotation of the LC molecules and perpendicular alignment of the molecules with respect to the waveguide (not to scale). [237]
@@
Various commercially-available LC compounds can be used for these modulators, with the compound used impacting parameters such as the birefringence at a given wavelength, the operating temperature range, and the modulation speed. For example, a commonly-utilized LC compound is 5CB, which has ordinary and extraordinary refractive indices of 1.53 and 1.7, respectively, at red wavelengths and a nematic-state operating temperature range of 18 to 35C.
@@
Integrated LC-based demonstrations have largely focused on operation at infrared wavelengths [235, 236, 238, 239, 240, 241, 242, 243, 244, 245]. Recently, integrated LC-based modulators have been shown at visible wavelengths by integrating nematic LC media into silicon-nitride-based platforms [237, 246, 247, 248, 249], enabling compact phase and amplitude modulation at a red operating wavelength. For example, an integrated LC-based phase modulator achieved $41\pi$ phase shift within ±$\pm 2.4$ V for a 500-$\upmu$m-long modulator, which means that a $2\pi$ phase shifter would need to be only 24.4 $\upmu$m long (corresponding to a 0.003 V-cm modulation requirement) [237, 246]. Moreover, an integrated LC-based amplitude modulator, which leverages the LC birefringence to tune the coupling coefficient between two waveguides, achieved amplitude modulation with a device length of only 14 $\upmu$[248, 247].
@@
Due to the large birefringence afforded by LC media, these integrated LC-based devices are one to two orders of magnitude more compact than comparable thermo-optic, electro-optic, or piezo-optomechanical modulators at visible wavelengths. While longer modulators may be adequate for certain systems and applications, many systems require compact modulators, such as integrated optical phased arrays that require dense integration for high fill factors. As such, these compact and power-efficient integrated LC-based modulators are well suited to enable densely-integrated visible-light systems for many impactful applications areas [250, 251, 7, 104, 8], including augmented-reality displays (as discussed in Sec. I). However, for applications requiring high-speed modulation, including most quantum and biological applications, the modulation speed of these integrated LC-based modulators will need to be improved.
III.6 Summary